# Q-Adapter: Customizing Pre-trained LLMs to New Preferences with Forgetting Mitigation

### v3

@@ Abstract
Large Language Models (LLMs), trained on a large amount of corpus, have demonstrated remarkable abilities. However, it may not be sufficient to directly apply open-source LLMs like Llama to certain real-world scenarios, since most of them are trained for general purposes. Thus, the demands for customizing publicly available LLMs emerge, but are currently under-studied. In this work, we consider customizing pre-trained LLMs with new human preferences. Specifically, the LLM should not only meet the new preference but also preserve its original capabilities after customization. Drawing inspiration from the observation that human preference can be expressed as a reward model, we propose to cast LLM customization as optimizing the sum of two reward functions, one of which (denoted as $r_{1}$) was used to pre-train the LLM while the other (denoted as $r_{2}$) characterizes the new human preference. The obstacle here is that both reward functions are unknown, making the application of modern reinforcement learning methods infeasible. Thanks to the residual Q-learning framework, we can restore the customized LLM with the pre-trained LLM and the residual Q-function without the reward function $r_{1}$. Moreover, we find that for a fixed pre-trained LLM, the reward function $r_{2}$ can be derived from the residual Q-function, enabling us to directly learn the residual Q-function from the new human preference data upon the Bradley-Terry model. We name our method Q-Adapter as it introduces an adapter module to approximate the residual Q-function for customizing the pre-trained LLM towards the new preference. Experiments based on the Llama-3.1 model on the DSP dataset and HH-RLHF dataset illustrate the superior effectiveness of Q-Adapter on both retaining existing knowledge and learning new preferences [^1] Code is available at https://github.com/mansicer/Q-Adapter.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in tasks such as natural language understanding (Wei et al., 2022), reasoning (Mondorf & Plank, 2024) and logic (Pan et al., 2023), and have already been successfully applied to diverse real-world scenarios including robotics (Zeng et al., 2023), healthcare assistants (Yuan et al., 2023) and scientific discovery (AI4Science & Quantum, 2023), to name a few. However, publicly available open-source LLMs, e.g., Llama 3 (Dubey et al., 2024) and Pythia (Biderman et al., 2023), frequently struggle in specialized domains which require expert knowledge (Wu et al., 2024) due to that they have been trained mainly on generic datasets. Further adaptation is therefore required to ensure that pre-trained LLMs excel in targeted use cases. Moreover, people often want to personalize the LLM by providing additional preferences and information (Kirk et al., 2024). How to effectively customize pre-trained LLMs with downstream data is thus one of the common demands for expanding the application scope of LLMs (Salemi et al., 2024; Wozniak et al., 2024; Liu et al., 2024).
Researchers and practitioners have explored various techniques to customize a pre-trained LLM to improve their performance on specific applications or domains. This includes methods like in-context learning (Gao & Das, 2024), task-specific fine-tuning (Liu et al., 2024), and retrieval augmented generation (Huang & Huang, 2024). In this paper, we consider customizing a pre-trained LLM with a task-specific preference dataset. Existing approaches to achieve this could be Supervised Fine-Tuning (SFT) on the preferred responses or Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022). However, the LLM may forget its capabilities after some rounds of fine-tuning (Luo et al., 2023; OpenAI, 2023), which is often not what we want. A natural question thus arises: How can we adapt the pre-trained LLM to downstream preferences while preserving its original knowledge and abilities, as demonstrated in Figure 1?
@@
In this paper, specifically, we focus our attention on LLMs that have been aligned with general human values via RLHF, like Llama (Dubey et al., 2024) and Gemma (Rivière et al., 2024) series, which we refer to as “pre-trained LLMs”. When it comes to customizing them, we notice that both the new human preference and general human values can be expressed as reward models (Ouyang et al., 2022), inspiring us to unify the aforementioned two requirements into one objective that maximizes the sum of two reward functions (see Section 3, one of which (denoted as $r_{1}$) models the general human values, while the other (denoted as $r_{2}$) characterizes the new human preference. Although both reward functions are unknown, this objective can be fulfilled by learning a residual Q-function without accessing the reward function $r_{1}$ (see Section 2.3 for a concise introduction to residual Q-learning (Li et al., 2023)). Moreover, we observe that given a fixed pre-trained LLM, the reward function $r_{2}$ and the residual Q-function are interchangeable, enabling us to directly learn the residual Q-function from the new human preference data upon the classical Bradley-Terry model (Bradley & Terry, 1952). Therefore, the customized LLM which learns the new human preference as well as inherits knowledge from the pre-trained LLM can be obtained by learning a residual Q-function without the reward function $r_{1}$ or $r_{2}$.
To summarize, our key contributions are as below:
@@
  • Our experiments on the Domain-Specific Preference (Cheng et al., 2023) dataset and the HH-RLHF (Bai et al., 2022) dataset show that Q-Adapter built upon Llama-3.1 models (Dubey et al., 2024) can effectively customize pre-trained LLMs to preserve existing knowledge and learn new preferences.
@@ 2 Preliminaries
In this section, we will first formalize the language generation task as a token-level Markov Decision Process (MDP) (Puterman, 1994), helping us to see how we can apply modern Reinforcement Learning (RL) approaches (Schulman et al., 2017; Haarnoja et al., 2018) to natural language processing problems. Then, we will present some basics of Reinforcement Learning from Human Feedback (RLHF). Finally, we will give a concise introduction to residual Q-learning (Li et al., 2023), which will serve as foundation for later derivation.
@@ 2.1 Language Generation as a Token-level MDP
Following Ramamurthy et al. (2023), we formalize the language generation task as an MDP, defined by a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{V},r,\mathcal{P},\gamma,\rho,T\rangle$, where $\mathcal{S}$ denotes the state space and the action space $\mathcal{V}$ is a finite vocabulary set. $\gamma\in[0,1]$ is a discount factor. Let $\pi:\mathcal{S}\to\Delta_{\mathcal{V}}$ be the LLM [^2] We use $\Delta_{X}$ to denote the set of distributions over $X$. (or policy). At the beginning of each episode in the MDP $\mathcal{M}$, a prompt $x=(x_{1},x_{2},\cdots,x_{m})$ of length $m$ is sampled from the initial state distribution $\rho$, with $m\in\mathbb{N}$ and $\forall i\in\quantity{1,2,\cdots,m},x_{i}\in\mathcal{V}$. We also call $x$ the initial state and denote it by $s_{1}\in\mathcal{S}$. At each time step $t\in\quantity{1,2,\cdots,T}$, the LLM selects an action (or equivalently, a token) $a_{t}\in\mathcal{V}$ according to $\pi(\cdot|s_{t})$. The environment then transits to the next state $s_{t+1}=(x,a_{1},\cdots,a_{t})$, rewarding the LLM with $r(s_{t},a_{t})\in R$. That is, the transition model $\mathcal{P}:\mathcal{S}\times\mathcal{V}\to\Delta_{\mathcal{S}}$ is deterministic. $\mathcal{P}(s_{t+1}|s_{t},a_{t})=1$ if and only if $s_{t+1}=s_{t}\oplus a_{t}$, where $\oplus$ means concatenation. The episode ends if an End-Of-Sentence (EOS) token is generated or the maximal length $T$ is reached. For notational simplicity, we assume that the length of any response is exactly $T$, with necessary padding occurring after the EOS token. The main objective in the MDP framework is to find an optimal policy which maximizes the expected cumulative sum of rewards.
@@ 2.2 Reinforcement Learning from Human Feedback
Learning the optimal LLM within the aforementioned MDP framework requires access to the ground-truth reward function, which can be infeasible for many language generation tasks. Classical RLHF methods, such as InstructGPT (Ouyang et al., 2022), thus considers first learning a reward model from human preferences and then optimizing it with RL algorithms. Given a dataset $\mathcal{D}=\quantity{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}}_{i=1}^{N}$ containing $N$ preference pairs, where $x^{(i)}$ represents the prompt, $y_{w}^{(i)}$ and $y_{l}^{(i)}$ are two corresponding responses with $y_{w}^{(i)}$ being the preferred one, we can train a reward model $r_{\phi}(s,a)$ with learnable parameters $\phi$ based on the Bradley-Terry model (Bradley & Terry, 1952), as demonstrated below:
@@
[…]
where $\alpha>0$ is a hyper-parameter, called entropy weight; $\pi_{\text{ref}}$ is a reference policy, which often results from SFT; $D_{\mathrm{KL}}$ is the Kullback-Leibler divergence (KL-divergence) (Kullback & Leibler, 1951), used for mitigating over-optimization of the reward model (Ouyang et al., 2022).
@@
$]Q^{\pi}(s,a)=\mathbb{E}\bqty{r_{\phi}^{\mathrm{KL}}(s,a)+\sum_{t=2}^{T}\gamma^{t-1}\pqty{r_{\phi}^{\text{KL}}(s_{t},a_{t})+\alpha\mathcal{H}\pqty{\pi(\cdot|s_{t})}}\Bigg{|}s_{1}=s,a_{1}=a}$
is the soft Q-function of the LLM $\pi$ with the expectation being taken over the randomness of $\pi$ and $\mathcal{P}$, i.e, $a_{t}\sim\pi(\cdot|s_{t})$, $)s_{t+1}\sim\mathcal{P}\pqty{\cdot|s_{t},a_{t}}$. The reward $r^{\text{KL}}_{\phi}(s_{t},a_{t})=r_{\phi}(s_{t},a_{t})+\alpha\log\pi_{\text{ref}}(a_{t}|s_{t})$.
@@ 2.3 Residual Q-learning
Let $\pi_{1}^{*}$ be a policy that was trained to be optimal with maximum entropy RL using a reward function $r_{1}:\mathcal{S}\times\mathcal{V}\to\mathbb{R}$ and entropy weight $\alpha_{1}$. $r_{2}:\mathcal{S}\times\mathcal{V}\to\mathbb{R}$ is another reward function. The situation we currently face may be that while we can obtain $\pi_{1}^{*}$, we are unable to obtain $r_{1}$. For example, many large models have been open-sourced on Hugging Face [^3] https://huggingface.co/, but the reward functions used to train these models have almost never been made available. The Residual Q-Learning (RQL) framework (Li et al., 2023) considers learning a policy $\tilde{\pi}^{*}$ to maximize the reward $\lambda r_{1}+r_{2}$ with the policy $\pi_{1}^{*}$ and the reward function $r_{2}$, given that the reward function $r_{1}$ is unknown. That is, $\tilde{\pi}^{*}\in\operatorname*{arg\,max}_{\pi}\sum_{t=1}^{T}\gamma^{t-1}\pqty{\lambda r_{1}(s_{t},a_{t})+r_{2}(s_{t},a_{t})-\tilde{\alpha}\log\pi(a_{t}|s_{t})}$ ), where $\lambda>0$ is a hyper-parameter controlling the weight of the reward function $r_{1}$. To get $\tilde{\pi}^{*}$ when $r_{1}$ is unknown, Li et al. (2023) define the residual Q-function as
$\hat{Q}(s,a)=\tilde{Q}^{*}(s,a)-\lambda Q_{1}^{*}(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{V},$
@@ 3 LLM Customization as Reward Maximization
Recall that in our setting, the LLM $\pi_{1}^{*}$ to be customized has been pre-trained with RLHF using an unknown reward function $r_{1}$. Thus to preserve the knowledge of $\pi_{1}^{*}$ after customization, the customized LLM $\tilde{\pi}^{*}$ should maximize the reward function $r_{1}$ as well. Moreover, suppose that another reward function $r_{2}$ can characterize the new human preference, then learning the new preference indicates that the customized LLM $\tilde{\pi}^{*}$ should also maximize the reward $r_{2}$ (Ouyang et al., 2022). This inspires us to propose the following objective:
@@ 4 Q-Adapter
Corollary 4.1.
Proof.
This concludes the proof. ∎
Corollary 4.1 indicates that for a fixed pre-trained LLM $\pi_{1}^{*}$, we can freely exchange between $r_{2}$ and $\hat{Q}$. Let $Q_{\theta}:\mathcal{S}\times\mathcal{V}\to\mathbb{R}$ be a function with learnable parameters $\theta$ (e.g., a learnable neural network) to approximate the residual Q-function $\hat{Q}$. The corresponding reward function $r_{2}$ can be computed as:

### v4

@@ Abstract
Large Language Models (LLMs), trained on a large amount of corpus, have demonstrated remarkable abilities. However, it may not be sufficient to directly apply open-source LLMs like Llama to certain real-world scenarios, since most of them are trained for general purposes. Thus, the demands for customizing publicly available LLMs emerge, but are currently under-studied. In this work, we consider customizing pre-trained LLMs with new human preferences. Specifically, the LLM should not only meet the new preference but also preserve its original capabilities after customization. Drawing inspiration from the observation that human preference can be expressed as a reward model, we propose to cast LLM customization as optimizing the sum of two reward functions, one of which (denoted as $r_{1}$) was used to pre-train the LLM while the other (denoted as $r_{2}$) characterizes the new human preference. The obstacle here is that both reward functions are unknown, making the application of modern reinforcement learning methods infeasible. Thanks to the residual Q-learning framework, we can restore the customized LLM with the pre-trained LLM and the residual Q-function without the reward function $r_{1}$. Moreover, we find that for a fixed pre-trained LLM, the reward function $r_{2}$ can be derived from the residual Q-function, enabling us to directly learn the residual Q-function from the new human preference data upon the Bradley-Terry model. We name our method Q-Adapter as it introduces an adapter module to approximate the residual Q-function for customizing the pre-trained LLM towards the new preference. Experiments based on the Llama-3.1 model on the DSP dataset and HH-RLHF dataset illustrate the superior effectiveness of Q-Adapter on both retaining existing knowledge and learning new preferences. Our code is available at https://github.com/LAMDA-RL/Q-Adapter.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in tasks such as natural language understanding (Wei et al., 2022), reasoning (Mondorf & Plank, 2024), and logic (Pan et al., 2023), and have already been successfully applied to diverse real-world scenarios including robotics (Zeng et al., 2023), healthcare assistants (Yuan et al., 2023b), and scientific discovery (AI4Science & Quantum, 2023), to name a few. However, publicly available open-source LLMs, e.g., Llama 3 (Team, 2024b) and Pythia (Biderman et al., 2023), frequently struggle in specialized domains which require expert knowledge (Wu et al., 2024) due to the fact that they have been trained mainly on generic datasets. Further adaptation is therefore required to ensure that pre-trained LLMs excel in targeted use cases. Moreover, people often want to personalize the LLM by providing additional preferences and information (Kirk et al., 2024). How to effectively customize pre-trained LLMs with downstream data is thus one of the common demands for expanding the application scope of LLMs (Salemi et al., 2024; Wozniak et al., 2024; Liu et al., 2024b).
Researchers and practitioners have explored various techniques to customize a pre-trained LLM to improve their performance on specific applications or domains. This includes methods like in-context learning (Gao & Das, 2024), task-specific fine-tuning (Liu et al., 2024b), and retrieval augmented generation (Huang & Huang, 2024). In this paper, we consider customizing a pre-trained LLM with a task-specific preference dataset. Existing approaches to achieve this could be Supervised Fine-Tuning (SFT) on the preferred responses or Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022). However, the LLM may forget its capabilities after some rounds of fine-tuning (Luo et al., 2023; OpenAI, 2023), which is often not what we want. A natural question thus arises: How can we adapt the pre-trained LLM to downstream preferences while preserving its original knowledge and abilities, as demonstrated in Figure 1?
@@
In this paper, specifically, we focus our attention on LLMs that have been aligned with general human values via RLHF, like Llama (Team, 2024b) and Gemma (Team, 2024a) series, which we refer to as “pre-trained LLMs”. When it comes to customizing them, we notice that both the new human preference and general human values can be expressed as reward models (Ouyang et al., 2022), inspiring us to unify the aforementioned two requirements into one objective that maximizes the sum of two reward functions (see Section 3), one of which (denoted as $r_{1}$) models the general human values, while the other (denoted as $r_{2}$) characterizes the new human preference. Although both reward functions are unknown, this objective can be fulfilled by learning a residual Q-function without accessing the reward function $r_{1}$ (see Section 2.3 for a concise introduction to residual Q-learning (Li et al., 2023)). Moreover, we observe that given a fixed pre-trained LLM, the reward function $r_{2}$ and the residual Q-function are interchangeable, enabling us to directly learn the residual Q-function from the new human preference data upon the classical Bradley-Terry model (Bradley & Terry, 1952). Therefore, the customized LLM which learns the new human preference as well as inherits knowledge from the pre-trained LLM can be obtained by learning a residual Q-function without the reward function $r_{1}$ or $r_{2}$.
To summarize, our key contributions are as follows:
@@
  • Our experiments on the Domain-Specific Preference (Cheng et al., 2023) dataset and the HH-RLHF (Bai et al., 2022) dataset show that Q-Adapter built upon Llama-3.1 models (Team, 2024b) can effectively customize pre-trained LLMs to preserve existing knowledge and learn new preferences.
@@ 2 Preliminaries
In this section, we will first formalize the language generation task as a token-level Markov Decision Process (MDP) (Puterman, 1994), helping us to see how we can apply modern Reinforcement Learning (RL) approaches (Schulman et al., 2017; Haarnoja et al., 2018) to natural language processing problems. Then, we will present some basics of Reinforcement Learning from Human Feedback (RLHF). Finally, we will give a concise introduction to residual Q-learning (Li et al., 2023), which will serve as a foundation for later derivation.
@@ 2.1 Language Generation as a Token-level MDP
Following Ramamurthy et al. (2023), we formalize the language generation task as an MDP, defined by a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{V},r,\mathcal{P},\gamma,\rho,T\rangle$, where $\mathcal{S}$ denotes the state space and the action space $\mathcal{V}$ is a finite vocabulary set. $\gamma\in[0,1]$ is a discount factor. Let $\pi:\mathcal{S}\to\Delta_{\mathcal{V}}$ be the LLM  [^1] We use $\Delta_{X}$ to denote the set of distributions over $X$. (or policy). At the beginning of each episode in the MDP $\mathcal{M}$, a prompt $x=(x_{1},x_{2},\cdots,x_{m})$ of length $m$ is sampled from the initial state distribution $\rho$, with $m\in\mathbb{N}$ and $\forall i\in\quantity{1,2,\cdots,m},x_{i}\in\mathcal{V}$. We also call $x$ the initial state and denote it by $s_{1}\in\mathcal{S}$. At each time step $t\in\quantity{1,2,\cdots,T}$, the LLM selects an action (or equivalently, a token) $a_{t}\in\mathcal{V}$ according to $\pi(\cdot|s_{t})$. The environment then transits to the next state $s_{t+1}=(x,a_{1},\cdots,a_{t})$, rewarding the LLM with $r(s_{t},a_{t})\in R$. That is, the transition model $\mathcal{P}:\mathcal{S}\times\mathcal{V}\to\Delta_{\mathcal{S}}$ is deterministic. $\mathcal{P}(s_{t+1}|s_{t},a_{t})=1$ if and only if $s_{t+1}=s_{t}\oplus a_{t}$, where $\oplus$ means concatenation. The episode ends if an End-Of-Sentence (EOS) token is generated or the maximal length $T$ is reached. For notational simplicity, we assume that the length of any response is exactly $T$, with necessary padding occurring after the EOS token. The main objective in the MDP framework is to find an optimal policy that maximizes the expected cumulative sum of rewards.
@@ 2.2 Reinforcement Learning from Human Feedback
Learning the optimal LLM within the aforementioned MDP framework requires access to the ground-truth reward function, which can be infeasible for many language generation tasks. Classical RLHF methods, such as InstructGPT (Ouyang et al., 2022), thus consider first learning a reward model from human preferences and then optimizing it with RL algorithms. Given a dataset $\mathcal{D}=\quantity{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}}_{i=1}^{N}$ containing $N$ preference pairs, where $x^{(i)}$ represents the prompt, $y_{w}^{(i)}$ and $y_{l}^{(i)}$ are two corresponding responses with $y_{w}^{(i)}$ being the preferred one, we can train a reward model $r_{\phi}(s,a)$ with learnable parameters $\phi$ based on the Bradley-Terry model (Bradley & Terry, 1952), as demonstrated below:
@@
$\max_{\pi}\mathbb{E}_{s_{1}\sim\rho,a_{t}\sim\pi(\cdot|s_{t})}\bqty{\sum_{t=1}^{T}\gamma^{t-1}\pqty{r_{\phi}(s_{t},a_{t})-\alpha D_{\mathrm{KL}}\pqty{\pi(\cdot|s_{t})\|\pi_{\mathrm{ref}}(\cdot|s_{t})}}},$ (2)
where $\alpha>0$ is a hyper-parameter, called entropy weight; $\pi_{\mathrm{ref}}$ is a reference policy, which often results from SFT; $D_{\mathrm{KL}}$ is the Kullback-Leibler divergence (KL-divergence) (Kullback & Leibler, 1951), used for mitigating over-optimization of the reward model (Ouyang et al., 2022).
@@
$))]Q^{\pi}(s_{1},a_{1})=\mathbb{E}\bqty{r_{\phi}^{\mathrm{KL}}(s_{1},a_{1})+\sum_{t=2}^{T}\gamma^{t-1}\pqty{r_{\phi}^{\mathrm{KL}}(s_{t},a_{t})+\alpha\mathcal{H}\pqty{\pi(\cdot|s_{t})}}}$ (4)
is the soft Q-function of the LLM $\pi$ with the expectation being taken over the randomness of $\pi$ and $\mathcal{P}$, i.e, $a_{t}\sim\pi(\cdot|s_{t})$, $)s_{t+1}\sim\mathcal{P}\pqty{\cdot|s_{t},a_{t}}$. The reward $r^{\mathrm{KL}}_{\phi}(s_{t},a_{t})=r_{\phi}(s_{t},a_{t})+\alpha\log\pi_{\mathrm{ref}}(a_{t}|s_{t})$.
@@ 2.3 Residual Q-learning
Let $\pi_{1}^{*}$ be a policy that was trained to be optimal with maximum entropy RL using a reward function $r_{1}:\mathcal{S}\times\mathcal{V}\to\mathbb{R}$ and entropy weight $\alpha_{1}$. $r_{2}:\mathcal{S}\times\mathcal{V}\to\mathbb{R}$ is another reward function. The situation we currently face may be that while we can obtain $\pi_{1}^{*}$, we are unable to obtain $r_{1}$. For example, many large models have been released on Hugging Face  [^2] https://huggingface.co/models, but the reward functions used to train these models have almost never been made available. The Residual Q-Learning (RQL) framework (Li et al., 2023) considers learning a policy $\tilde{\pi}^{*}$ to maximize the expected discounted sum of the reward $\lambda r_{1}+r_{2}$ with the policy $\pi_{1}^{*}$ and the reward function $r_{2}$, given that the reward function $r_{1}$ is unknown. That is, $))]\tilde{\pi}^{*}\in\operatorname*{arg\,max}_{\pi}\mathbb{E}_{s_{1}\sim\rho}\bqty{\sum_{t=1}^{T}\gamma^{t-1}\pqty{\lambda r_{1}(s_{t},a_{t})+r_{2}(s_{t},a_{t})+\tilde{\alpha}\mathcal{H}\pqty{\pi(\cdot|s_{t})}}}$, where $\tilde{\alpha}$ is the entropy weight and $\lambda>0$ is a hyper-parameter controlling the importance of the reward function $r_{1}$. To get $\tilde{\pi}^{*}$ when $r_{1}$ is unknown, Li et al. (2023) define the residual Q-function as
$\hat{Q}(s,a)=\tilde{Q}^{*}(s,a)-\lambda Q_{1}^{*}(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{V},$ (5)
@@ 3 LLM Customization as Reward Maximization
Recall that in our setting, the LLM $\pi_{1}^{*}$ to be customized has been pre-trained with RLHF using an unknown reward function $r_{1}$. Thus to preserve the knowledge of $\pi_{1}^{*}$ after customization, the customized LLM $\tilde{\pi}^{*}$ should maximize the expected discounted sum of the reward $r_{1}$ as well. Moreover, suppose that another reward function $r_{2}$ can characterize the new human preference, then learning the new preference indicates that the customized LLM $\tilde{\pi}^{*}$ should also maximize the expected discounted sum of the reward $r_{2}$ (Ouyang et al., 2022). This inspires us to propose the following objective:
@@ 4 Q-Adapter
Corollary 4.1.
Proof.
This concludes the proof. ∎
Corollary 4.1 indicates that for a fixed pre-trained LLM $\pi_{1}^{*}$, we can freely exchange between $r_{2}$ and $\hat{Q}$. Let $Q_{\theta}:\mathcal{S}\times\mathcal{V}\to\mathbb{R}$ be a function with learnable parameters $\theta$ (e.g., a learnable neural network) to approximate the residual Q-function $\hat{Q}$. That is, we can compute the reward function $r_{2}$ via